UNIT 5 – QUESTION 2 (25 marks) Given the following information: Demand: Qd = 200 – 5P Supply: Qs = 5P If a quantity tax of $2 per unit sold is imposed,

(a)Considering that the government will earn revenue, overall, do you think that the society benefits from the imposition of the tax? Explain.

(b) Calculate the equilibrium market price and the equilibrium quantity sold.

(c) Determine the demand and supply equation after the tax.

(d) What will be the new equilibrium price paid by the buyers and the new price received by the supplier?

(e) Calculate the new equilibrium quantity sold. (f) Calculate the tax revenue earned by the government.

(g) Calculate the deadweight loss due to the tax. (h) What determines whether the buyer of the seller bears the burden of the tax?
Answer #1

1. (a) gf goverment will earn revenue from tax then society does not besitit from imposstrom of tax. because face impositionStr817 (e) 10) fae pps unit Q!= 5/p-2) - SP-10 =)@=-10+5P1 @ = 200-51 la Demand qt marret Equilibrium after tax Qo Qaraw 200